Log5x-Log(2x-1)=Log7
Answer by lwsshak3(11628)

Log5x-Log(2x-1)=Log7
log[5x/(2x-1)]=log(7)
5x/(2x-1)=7
5x=14x-7
9x=7
x=7/9